OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x2 + 6x + 7 = 0

OpenStudy (anonymous):

To complete the square, use the fact that x^2+bx=(x+0.5b)^2-0.25b^2

OpenStudy (anonymous):

Now re-write the equation as (x^2+6x)+7=0 and you would notice that b=6

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

Alright.... x^2+6x+7=0 (x^2+6x)+7=0 [(x+3)^2-9]+7=0 (x+3)^2-2=0 (x+3)^2=2

OpenStudy (anonymous):

"Completing the square" operation is simply, whenever you see this term x^2+bx (for any constant b), you can re-write it as x^2+bx=[x+(b/2)]^2-(b^2/4)

OpenStudy (anonymous):

here you have the term x^2+6x which can be rewritten as (x+3)^2-9

OpenStudy (anonymous):

Go here the answer is their! :) http://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.7474.html

OpenStudy (anonymous):

Answer ---> (x+3)^2-16=0

OpenStudy (anonymous):